OpenStudy (anonymous):

Is anyone good at solving systems of equations with three variables? 6x + 6y + 5z = -21 3x - y + 6z = 20 -3x +4y +2z = -11 Mahalo nui loa~

OpenStudy (chaise):

Do you want the answer(s) or do you want to know how to solve equation like these?

OpenStudy (anonymous):

I just want to know how to solve an equation like these. The problem was just an example.

OpenStudy (chaise):

Okay - I hope you're familiar with solving equations with two variables and two equations? 6x + 6y + 5z = -21 3x - y + 6z = 20 -3x +4y +2z = -11 6x + 6y +5z =-21 3x -y + 6z = 20<-- multiply by two 6x + 6y +5z =-21 6x-2y+12z = 40 Eliminate by subtration. 6y-7z=-61 3x - y + 6z = 20 -3x +4y +2z = -11 Eliminate by addition. 3y + 8z = 9 6y-7z = -61 3y + 8z = 9 Do you know how to do the rest?

OpenStudy (anonymous):

Yes, I know how to do the rest. Thank you! :D

OpenStudy (chaise):

Just make sure you solve for each y, z and x, and not just your y and z. :) Goodluck!